OpenStudy (anonymous):

solve each equation 3x^3+12x^2-x-4=0

OpenStudy (anonymous):

I made a mistake putting it up! just edited. can you look again please

OpenStudy (anonymous):

3x^3+12x^2-x-4=0 3x^2(x+4)-(x+4)=0 which is same as 3x^2(x+4)+(-1)(x+4)=0 (3x^2-1)(x+4)=0 3x^2-1=0 3x^2=1 x^2=1/3 x= +-1/3 and x+4=0 x=-4